# SSCATeR: Sparse Scatter-Based Convolution Algorithm with Temporal Data Recycling for Real-Time 3D Object Detection in LiDAR Point Clouds

### v1

@@ Abstract
{IEEEkeywords} Data Reuse, Object Detection, Scatter Operation, Sparse Convolution, Streaming LiDAR.
@@ 2.3 LiDAR as a Streaming Sensor
Sparse PointPillars [sparse_pointpillars] Use sparse convolutions to reduce the inference time of PointPillars for real-time embedded processing. Still utilizes 100ms100\ ms frames, meaning that real-time processing in their use-case still creates a 200ms200\ ms delay. Also follows a sparse approach, but utilizes temporally sequential frames and data reuse to further reduce computation.
PillarNet [pillarnet] Use sparse convolutions and a powerful feature encoder to address the gap between pillar- and voxel-based point cloud object detection Still utilizes 100ms100\ ms frames, meaning that real-time processing in their use-case still creates a 162.5ms162.5\ ms delay.
@@ 3.1 Dataset Collection
To enable model training and testing, we collected multi-drone survey datasets at two contrasting locations in Ireland and Germany. The first location is a railway bridge in Newcastle, County Wicklow, Ireland [denoise]. Two DJI M300 drones are utilized as shown in Fig. 1. One drone carried the DJI Zenmuse L1 LiDAR system as its payload for data collection. The six channel L1 LiDAR utilizes a non-repetitive scanning pattern as shown in Fig. 2 for a $70.4\degree\times 77.2\degree$ field-of-view (FoV) with a pulse rate of 240,000 points per second. The reconstruction of the collected survey point cloud is displayed in Fig. 3. The DJI L1 system, however, cannot be used for real-time in-field testing, as the data is not accessible in real-time. As such, we have created an onboard compute system which utilizes the Livox Avia, the LiDAR unit found inside the L1, which is pictured in Fig. 4.
@@ 3.3 Sparse Scatter-Based Convolutions with Temporal Data Recycling: SSCATeR
Figure 11: A time series of pseudo-images in 10ms10\ ms increments over a 100ms100\ ms interval. In this example, the pillars brightness represents the maximum intensity of the points in the pillar.
Figure 12: The 10ms10\ ms increment binary change maps of the pseudo-images seen in Fig. 11.
@@ 4.4 Comparison of Different Architectures
Architecture Precision Recall F1-Score mAP (IoU @ 0.3)
@@ 4.6 Performance and Results Evaluation
There are failure cases for the SSCATeR algorithm, for example, when the LiDAR scanning pattern does not capture points belonging to the drone due to its non-repetitive nature. As shown in Fig. 25, a missed detection can propagate across multiple frames. The use of a tracking algorithm could provide a fallback sub-system in the case of a full SAD system.

### v2

@@ Abstract
{IEEEkeywords} Point Cloud, Object Detection, Scatter Operation, Sparse Convolution, Streaming LiDAR, PointPillars, Crosstalk Interference.
@@ 2.3 LiDAR as a Streaming Sensor
SpikeClouds [spikeclouds] is a streaming LiDAR object detection network which utilizes a combination of a spiking neural network (SNN) backbone and a CNN-based neck and detection head. SpikeClouds uses this architecture to process individual scanlines produced by MEMS-based LiDARs, in contrast to previous works which focus on rotational scanners. The SNN processes each scanline individually and the CNN processes the combined SNN output once per scan. As such, backbone processing occurs throughout the acquisition time, and only neck and detection head processing have to wait for the full scan completion. Their work, therefore, still requires a full 100ms100\ ms scan before processing, leading to an end-to-end processing time of 120ms120\ ms on an Nvidia 2080 GPU. In terms of performance, SpikeClouds achieves an mAP score within $10$ points of PillarNet, and is a notable contribution which broadens LiDAR object detection away from work which focuses solely on rotational scanners.
@@
Sparse PointPillars [sparse_pointpillars] Use sparse convolutions to reduce the inference time of PointPillars for real-time embedded processing. Utilizes 100ms100\ ms frames, meaning that real-time processing in their use-case still creates a 200ms200\ ms delay. Also follows a sparse approach, but utilizes temporally sequential frames and data reuse to further reduce computation.
PillarNet [pillarnet] Use sparse convolutions and a powerful feature encoder to address the gap between pillar- and voxel-based point cloud object detection Utilizes 100ms100\ ms frames, meaning that real-time processing in their use-case still creates a 162.5ms162.5\ ms delay.
@@
SpikeClouds [spikeclouds] Processes LiDAR scanlines through an SNN backbone during acquisition, meaning only neck and detection head processing must wait for the full LiDAR scan before processing. Is able to work with non-rotational LiDAR scanners. Requires a full LiDAR scan before detections can be processed. Does not require a full LiDAR scan to complete detections. Also works with non-rotational LiDAR scanners.
@@ 3.1 Dataset Collection
To enable model training and testing, we collected multi-drone survey datasets at two contrasting locations in Ireland and Germany. The first location is a railway bridge in Newcastle, County Wicklow, Ireland [denoise]. Two DJI M300 drones are utilized as shown in Fig. 1. One drone carried the DJI Zenmuse L1 LiDAR system as its payload for data collection. The five channel L1 LiDAR utilizes a non-repetitive scanning pattern enabled by the use of Risley prisms as shown in Fig. 2 for a $70.4\degree\times 77.2\degree$ field-of-view (FoV) with a pulse rate of $240,000$ points per second. The reconstruction of the collected survey point cloud is displayed in Fig. 3. The DJI L1 system, however, cannot be used for real-time in-field testing, as the data is not accessible in real-time. As such, we have created an onboard compute system which utilizes the Livox Avia, the LiDAR unit found inside the L1, which is pictured in Fig. 4.
@@ 3.3 Sparse Scatter-Based Convolutions with Temporal Data Recycling: SSCATeR
Figure 11: A time series of pseudo-images in 10ms10\ ms increments over a 100ms100\ ms interval. In this example, the pillars brightness represents the maximum intensity of the points in the pillar. The location of the drone is marked using a green box in each 10ms10\ ms increment.
Figure 12: The 10ms10\ ms increment binary change maps of the pseudo-images seen in Fig. 11. The location of the drone is marked using a green box in each 10ms10\ ms increment.
@@ 3.6 PandaSet Testing
3.7 LiDAR-to-LiDAR Crosstalk Interference Testing
Given the use case of multi-drone swarms, we also explore the effects of crosstalk interference between two drone-mounted LiDAR units on the accuracy of the network. To test this, we have collected a dataset at Strathaven Airfield, using two L1 LiDAR systems mounted on two DJI M300 drones during a survey of an aircraft hangar. We have included scenarios of both indirect crosstalk and direct crosstalk. For the indirect crosstalk, where pulses from one LiDAR may reflect off a surface and erroneously be captured by another LiDAR, we stationed the drones at $45\degree$ and $45\degree$ angles facing toward the hangar, as well as having the surveying drone move directly in front of the monitoring drone. For the direct crosstalk, where pulses from one LiDAR are fired towards and enter another LiDAR system directly, the two drones faced one another with their LiDARs facing forward. We utilize $3,635$ frames for training, and $1,810$ frames for testing. We train both models over $128$ epochs using the same hyperparameters as those described in section 3.5.
@@ 4.4 Comparison of Different Architectures
Architecture Precision Recall F1-Score mAP
@@ 4.5 Testing on PandaSet
4.6 LiDAR-to-LiDAR Crosstalk Interference Results
Table 12:
Architecture Crosstalk Type Precision Recall F1-Score mAP
PillarNet (0.16m0.16\ m) Baseline $94.192$ $94.972$ $94.580$ $93.855$
No Indirect $94.866$ $94.972$ $94.920$ $93.877$
No Direct $94.554$ $94.972$ $94.763$ $93.870$
No Crosstalk $95.235$ $94.972$ $95.103$ $93.892$
Aerial-PointPillars Baseline $83.588$ $93.702$ $88.356$ $93.021$
No Indirect $84.504$ $93.702$ $88.866$ $93.051$
No Direct $84.085$ $93.702$ $88.633$ $93.029$
No Crosstalk $85.013$ $93.702$ $89.146$ $93.059$
This section discusses and analyzes the results of the LiDAR-to-LiDAR crosstalk interference testing, as described in section 3.7. We measure the performance of the PillarNet and Aerial-PointPillars models across a baseline, no direct crosstalk interference, no indirect crosstalk interference, and no crosstalk interference of either type. To categorize false positive predictions as attributable to crosstalk, visual analysis of prediction boxes was conducted, and those found to be containing or directly adjacent to crosstalk points were classified as such. Our testing found no false negatives that could be reasonably attributed to crosstalk. The results of this testing can be found in Table 12. For PillarNet, of the $106$ false positives, $20$ were attributed to crosstalk, with seven being caused by direct interference, and twelve from indirect interference. The mAP of the network increases by $0.037$ points, from $93.855\%$ to $93.892\%$. For Aerial-PointPillars, of the $332$ false positives, $33$ were attributed to crosstalk, with twelve being caused by direct interference, and $21$ from indirect interference. The mAP of the network increases by $0.038$ points, from $93.021\%$ to $93.059\%$. An example of a false positive created by crosstalk can be seen in Fig. 23.
Figure 23: Example of a false positive caused by indirect LiDAR-to-LiDAR crosstalk interference, detected using the trained Aerial-PointPillars model.
Of the $541$ test frames collected whilst the LiDAR’s were directly facing each other, false positives occurred in $1.29\%$ of inference passes for PillarNet and $2.21\%$ for Aerial-PointPillars. Of the $1,269$ frames collected during scenarios where indirect crosstalk was likely to occur, false positives occurred in $1.02\%$ of inference passes for PillarNet and $1.65\%$ for Aerial-PointPillars. Analysis of the point clouds shows that, while direct crosstalk did create much more interference, the intensity of the points was much higher than that of the drone points the network has been trained on, whereas indirect crosstalk intensity was similar to those belonging to a drone, and we believe this to be the cause of the similarity in the occurrence of false positives.
@@ 4.8 Analysis and Discussion of Method Limitations
There are failure cases for the SSCATeR algorithm, for example, when the LiDAR scanning pattern does not capture points belonging to the drone due to its non-repetitive nature. As shown in Fig. 26, a missed detection can propagate across multiple frames. Additionally, if too few points are collected in the leading 10ms10\ ms window to make a detection, it may be required that subsequent 10ms10\ ms frames are required to build up the feature map delaying the time of achieving successful detection. The use of a tracking algorithm could provide a fallback sub-system in the case of a full SAD system. From a latency perspective, SSCATeR performs best when point clouds are extremely sparse, denser point clouds can limit its capability to perform in real-time, and therefore much consideration is required in hardware design, not just in terms of SWaP restrictions and compute power, but additionally for parameters such as pulses per second from the LiDAR unit itself.
@@
Considering the results of the crosstalk testing, we find the interference to be minimally detrimental to object detection performance. Crucially from a safety critical aspect, crosstalk did not affect the false negatives and thus the recall, meaning that the only errors were in the form of additional detections rather than missed detections. That being said, there is still the potential that crosstalk could affect the morphology of a drone’s point cloud, and thus impact the recall, as such mission planning with multiple LiDARs must minimize the effects of crosstalk interference. Reducing direct interference is rather straightforward, drones in a swarm should not directly face one another with their LiDAR sensors during operation. The reduction of indirect interference is more complicated and should consider the beamforming of the LiDAR sensor. In our testing, the majority of indirect crosstalk occurred when the center of two sensors were aligned and facing the same direction. Therefore, ensuring that the observer drone is located at a different altitude to the LiDAR drones within the swarm should help reduce the effects of indirect crosstalk. Furthermore, the number of erroneous points caused by the crosstalk interference is limited by the design of the chosen LiDAR units. In Risley prism-based LiDAR units, the emitter and the photodetector used to receive the return sit behind two rotating prisms, meaning that crosstalk interference must enter the prisms at the correct angle to reach the photodetectors. This further supports the use of non-repetitive scanning patterns when using LiDARs in drone swarm-based object detection.